OpenStudy (anonymous):

How to Evaluate the following expressions and what we call these type of expressions?

OpenStudy (anonymous):

\[8^{1/3}\]\[\sqrt{27}/\sqrt{27}\]

OpenStudy (anonymous):

I think I should attach the picture so you guys see much better. Please tell me how do I solve these type of expressions and what we call these expressions?

OpenStudy (anonymous):

do you know what 4^(1/2) is?

OpenStudy (anonymous):

No, can you please explain it?

OpenStudy (anonymous):

its the same as sqrt(4) do you know how to do square roots?

OpenStudy (anonymous):

No, please explain little more. What do we call these type of expressions in algebra? I have found this exercise on Wikipedia that's why I am asking. I have no idea about how to take square roots of numbers. Please if you explain it.

OpenStudy (anonymous):

what is 2 times 2 what is 3 times 3 what is 4 times 4 tell me these things and il show you square roots

OpenStudy (anonymous):

4 9 16

OpenStudy (anonymous):

the square root of 4 is 2 the square root of 9 is 3 the square root of 16 is 4 you get what im saying

OpenStudy (anonymous):

Yes, I got it. It means the opposite of square

OpenStudy (anonymous):

so what is the cube root of 8

OpenStudy (anonymous):

cube root of 8 is 512

OpenStudy (anonymous):

no thats 8 cubed

OpenStudy (anonymous):

No, it is 2 ? right

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

theres your first expression

OpenStudy (anonymous):

but how do I know that I have to take the cube root or square root of the quantity while it looks like this on the Exponential power 1/4 ?

OpenStudy (anonymous):

so the cube root of 8 is 2 becuase 2*2*2 = 8. Cube = times itself 3 times 1/4th root is times itself 4 times. For example the square root of 16 is 4, but the 1/4th root of 16 is 2 In the same way taking 8^(2/3) is the same as taking it 8^2 and 8^(1/3) in any particulare order.

OpenStudy (anonymous):

so you can take 16^(1/4) = 2 becuase 2*2*2*2 = 16, 32^(1/5) is 2 becuase 2*2*2*2*2 is 32

OpenStudy (anonymous):

So, it means we have to take the root of the denominator of any expression?

OpenStudy (anonymous):

I mean denominator of Exponential power

OpenStudy (anonymous):

yes, if it is a fraction you take teh root, if it is a number you just take it times itself that many times, the order is unimportant, as long as you do them all

OpenStudy (anonymous):

also sqrt(1/2) = 1/sqrt(2)

OpenStudy (anonymous):

not all square roots turn out to be nice numbers, if they dont you usually just leave it in square root form

OpenStudy (anonymous):

terrible drawing but you get the point

OpenStudy (anonymous):

(1/8)^1/3 = 1/8?

OpenStudy (anonymous):

(1/8)^(1/3) = 1/[8^(1/3)]

OpenStudy (anonymous):

you already know 8^(1/3)

OpenStudy (anonymous):

also anything to the negative number just means you do 1/(that number) and make it positive again n^(-1) = 1/n 1/n^(-1) = n

OpenStudy (anonymous):

anything to the zero = 1

OpenStudy (anonymous):

Thanks, now how do we solve this type of expression? \[\sqrt{27} /\sqrt[3]{9}\]

OpenStudy (anonymous):

anything to 1 = itself

OpenStudy (anonymous):

well there is a trick you can use, but ultimately you wont get a nice number becuase sqrt(27) is not a nice number however you can draw out numbers from squares for example 8^(1/2) = 4^(1/2) * 2^(1/2) = 2 * 2^(1/2)

OpenStudy (anonymous):

what can you write sqrt(27) as to make it simpler

OpenStudy (anonymous):

you understand how I took 4 out of 8 right?

OpenStudy (anonymous):

3^3

OpenStudy (anonymous):

well yea 27 is 3^3 but sqrt(27) is actually 3^(3/2)

OpenStudy (anonymous):